40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 9 11:41:09 2015, ericq, Update, VAC, N2 Pressure fell 
    Reply  Fri Oct 16 18:27:52 2015, Steve, Update, VAC, vacuum normal condition oct16Fpm2015.png
       Reply  Tue Oct 20 16:27:04 2015, ericq, Update, VAC, vacuum VME machines rebooted 
          Reply  Wed Oct 21 10:03:15 2015, Steve, Update, VAC, after running out of N2 1d_after_reboot.png
             Reply  Wed Oct 21 15:42:56 2015, Steve, Update, VAC, vacuum gauge check Gauges.png
                Reply  Tue Oct 27 10:08:12 2015, Steve, Update, VAC, CC vacuum gauges gauges_5y_vs_8d.png13.9yCC1&4.png
                   Reply  Mon Dec 7 11:17:51 2015, Steve, Update, VAC, noisy RGA scan at day 433 433dRGAscan.png
                      Reply  Tue Dec 8 15:40:48 2015, Steve, Update, VAC, glitchy RGA scan at day 434 lessGlichingToday.png
                         Reply  Wed Dec 9 11:25:55 2015, Steve, Update, VAC, normal RGA scan at day 435-436 glichesGone.pngd436.png
                            Reply  Fri Jan 29 09:59:23 2016, Steve, Update, VAC, vacuum control upgrade 1100days.png
             Reply  Thu Oct 22 08:52:04 2015, Steve, Update, VAC, clean RGA scan after sweaty Maglev clean_scan.png
          Reply  Fri Sep 30 10:37:57 2016, Steve, Update, VAC, vacuum VME machines rebooted rebooted.png
Message ID: 11706     Entry time: Wed Oct 21 15:42:56 2015     In reply to: 11705     Reply to this: 11719
Author: Steve 
Type: Update 
Category: VAC 
Subject: vacuum gauge check 

Convectron gauge check:

Brand new 10 years old convectron gauge at atm was swapped into the place of existing gauges to see if they read close to 760 Torrs

They did reasonable well at the low end. I tried to imitate calibration and bring down the high end with little success.

( P1 and P2  were reading 7e-4 to ~660 Torr  The correction of the upper end pushed up the lower end too. I will correct this later

P3 and P4 high ends are way off )

The point is that they work.  Convectron gauges will be replaced and calibrated at the next vent.

Interlocks were not triggered during this test. I was expected to close the PSL shutter when P1 was reading 760 Torr

This hide some problem or  not understanding.

It was good to see CC1 and CC4 working at the moment

 

Somehow I succeded opening VM1 and closed VM2 = vacuum normal

I just hope it stays open overnight to get comparable RGA scan.

 

Attachment 1: Gauges.png  117 kB  | Hide | Hide all
Gauges.png
ELOG V3.1.3-